OpenStudy (anonymous):

A hemispherical container with a radius of 2m. When the depth of the liquid is h metres, the volume of the liquid, V m^3 is given: V= [ pih^2(6-h) ]/3 If the liquid is being pumped out at a constant rate of 0.25m^3/min, find the rate at which h is falling at the instant that h=1. (answer to nearest mm/minute) Also, prove that the given formula for V to be correct.

OpenStudy (anonymous):

Integral of rotation is a bit tedious, but doable

OpenStudy (anonymous):

Look I'll give you an outline @Kystal : You have a function\[V(h) = K* (6 h^2 -h^3)]\]

OpenStudy (anonymous):

this gives you the derivative

OpenStudy (anonymous):

\[\frac{dV}{dh}\]

OpenStudy (anonymous):

Now using the "inverse function derivative" theorem compute\[\frac{ dh }{ dV }\]

OpenStudy (anonymous):

Now the rate of flow is in fact the rate of change of VOLUME\[\frac{ dV }{ dt }\]

OpenStudy (anonymous):

erm... why is it K∗(6h2−h3)]? where the K comes from?

OpenStudy (anonymous):

Multiply and use the chain rule\[\frac{dh}{dV}*\frac{dV}{dt} = \frac{dh}{dt}\]

OpenStudy (anonymous):

\[K = \frac{\pi}{6}\]

OpenStudy (anonymous):

sorry pi/3

OpenStudy (anonymous):

This is a 35 minutes (at min) work

OpenStudy (anonymous):

That counts the proof of volume formula

OpenStudy (anonymous):

Is there a picture or diagram with this?

OpenStudy (anonymous):